# The dimension and Bose distance of certain primitive BCH codes

### v2

@@ II Preliminaries
with $\mathcal{G}=\bigcup\limits_{a=b}^{b+\delta-2}C_{a}.$ It is clear that $\mathrm{deg}(g(x))=\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{a}\right|$, where $\mathrm{deg}$ denotes the degree of a polynomial, and ||$\cdot$|| ⋅ | denotes the size of a set. The dimension of $\mathcal{C}_{(q,m,\delta)}$ can be given by
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta,b)})=n-\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{a}\right|=n-\sum\limits_{a\in\mathcal{L}(b,\delta)}\left|C_{a}\right|,$ (2)
where $\mathcal{L}(b,\delta)$ denotes the set of coset leaders in $[b,b+\delta-2].$ Additionally, for $\delta^{{}^{\prime}}\geq\delta$, we can observe that
$\bigcup\limits_{a=b}^{b+\delta-2}C_{a}=\bigcup\limits_{a=b}^{b+\delta^{{}^{\prime}}-2}C_{a}$
if and only if any integer in $[b+\delta-1,b+\delta^{{}^{\prime}}-2]$ is not a coset leader. Consequently, by recalling the definition of Bose distance $d_{B}$, we have
$d_{B}(\mathcal{C}_{(q,m,\delta,b)})=\delta^{{}^{\prime}}+1-b,$ (3)
where $\delta^{{}^{\prime}}$ denotes the smallest coset leader in $[b+\delta-1,n-1]$. In particular, the Bose distance of $\mathcal{C}_{(q,m,\delta)}$ is equal to the smallest coset leader in $[\delta,n-1]$. Therefore, we have
$d_{B}(\mathcal{C}_{(q,m,\delta,b)})=d_{B}(\mathcal{C}_{(q,m,b+\delta-1)})+1-b.$ (4)
For simplicity, we use $a\mid b$ to denote that the integer $a$ divides the integer $b$, and $a\nmid b$ to denote that $a$ does not divide $b$. It is well known that a necessary condition for an integer $a$ being the coset leader of $C_{a}$ is $q\nmid a$. Therefore, we only need to focus on finding those integers that are neither divisible by $q$ nor a coset leader. We define $S$ as the set of all such integers in $[1,n-1]$. That is,
$\mathcal{S}$ $=$ $\{a\in[1,n-1]:q\nmid a\hbox{ and }a\hbox{ is not a coset leader}\}.$ (5)
@@
[47, Theorem 2.3] When $m$ is an odd integer, for any integer $a\leq q^{h+1}$, $a$ is the coset leader of $C_{a}$ if and only if $q\nmid a$. When $m$ is an even integer, for any integer $a\leq 2q^{h}$, $a$ is the coset leader of $C_{a}$ if and only if $q\nmid a$.
III New results on $q$-cyclotomic cosets
@@
When $m$ is an even integer, for any integer $a\in[1,q^{m-\lfloor{m}/{4}\rfloor}),$
@@
where $\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell}$ forms the $q$-adic expansion of the integer $a\in[0,n].$ Let $I_{m-1}$ be the identity matrix of order $m-1$, and let
@@
When $m$ is an even integer, we can use a similar argument as in the first paragraph to show that $l_{a}>\lfloor{m}/{4}\rfloor$ for any integer $a\in[1,q^{m-\lfloor{m}/{4}\rfloor}).$ Note that $l_{a}$ divides $m$ and $m$ is even. It follows that $l_{a}={m}/{2}$ or $m$. Consequently, we have $l_{a}={m}/{2}=h$ if $aq^{h}\mathrm{\ mod\ }n=a$, and $l_{a}=m$ otherwise. Equivalently, the equality in (8) holds. This completes the proof. ∎
@@
Let $\mathcal{H}$ the set of all integers $a\in[1,n-1]$ such that $a$ is a coset leader and $|C_{a}|={m}/{2}$. That is,
$\mathcal{H}$ $=$ $\{a\in[1,n-1]:a\hbox{ is a coset leader and }|C_{a}|={m}/{2}\}.$ (11)
@@
$(\mathbf{0}_{h-k-1},a_{h+k}\ldots,a_{0})=(a_{h-1},\ldots,a_{0},\mathbf{0}_{h-k-1},a_{h+k}\ldots,a_{h}).$
@@
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}\cdot q^{2i-1}\quad\hbox{for }i\in[1,k],$
and
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}+\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}$ $=$ $\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}\cdot q^{2i-1}\quad\hbox{for }i\in[-k+1,0].$
Noticing that $10\leq\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}<$ 1 for $i\in[-k+1,0]$, and the two summations in (19) are integers, we can conclude that (19) is equivalent to
@@
If $a\in\mathcal{A}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell},$ then $i$ is the smallest integer in $[-k+1,k]$ such that $a_{h+i}>0$. Consequently, $\mathcal{A}_{k}(i)\cap\mathcal{A}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k+1,k]$.
Similarly, if $a\in\mathcal{B}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$, then $i$ is the smallest integer in $[-k,k]$ such that $a_{h+i}>0$. As a result, $\mathcal{B}_{k}(i)\cap\mathcal{B}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k,k]$.
@@
$V(a)Q^{h-i+1}=(\mathbf{0}_{h-k},a_{k+i-1},\ldots,a_{0},\mathbf{0}_{h-k},a_{h+k},\ldots,a_{h+i}).$ (25)
@@
With (15), it follows that $(a_{k+i-1},\ldots,a_{2i-1})<(a_{h+k},\ldots,a_{h+i}).$ Consequently, we still have $V(a)Q^{h-i+1}<V(a)$, which implies $a\in\mathcal{S}$. By now, we have demonstrated that $\bigcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i)\subseteq\mathcal{S}\cap[q^{h+k},q^{h+k+1}).$
Conversely, let us assume that $a\in\mathcal{S}\cap[q^{h+k},q^{h+k+1})$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$. We first have $a_{h+k}>0$. By the definition of $\mathcal{S}$ and Observation 3, we also have $a_{0}>0$ and $V(a)Q^{t}<V(a)$ for some integer $t\in[1,m-1]$. Noticing that $V(a)=(\mathbf{0}_{h-k},a_{h+k},\ldots,a_{0})$, it follows that the first $h-k$ entries of $V(a)Q^{t}$ are all equal to zero. Considering $a_{h+k}>0$ and $a_{0}>0$, we have $h-k+1\leq t\leq h+k$. Consequently, we can conclude that
@@
Given that $a_{h+k}>0$, we can chose $i$ to be the smallest integer in $[m-h-t,k]$ such that $a_{h+i}>0$. Then $V(a)$ has the form specified in (16) with $(a_{k+i-1},\ldots,a_{0})=(\mathbb{0}_{h+t+i-m},a_{h+k-t},\ldots,a_{0})$. Furthermore, the inequality $h-k+1\leq t\leq h+k$ implies that $i\in[-k+1,k]$.
Additionally, if $i>m-h-t$, we can obtain (15) from the fact that $a_{h+k}>0$ and $a_{0}>0$. On the other hand, if $i=m-h-t$, Then we can derive (15) from $V(a)Q^{t}<V(a)$ and $a_{0}>0$. Now we can conclude that $a\in\mathcal{A}_{k}(i)$ for some integer $i\in[-k+1,k]$. It follows that $\mathcal{S}\cap[q^{h+k},q^{h+k+1})\subseteq\bigcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i).$
We can conclude from the above argument that $\bigcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i)=\mathcal{S}\cap[q^{h+k},q^{h+k+1}).$ Finally, applying Remark 6 directly yields equation (23).
@@
IV The dimension of $\mathcal{C}_{(q,m,\delta)}$ for $2\leq\delta\leq q^{\lfloor({2m-1})/{3}\rfloor+1}$
Let $\mathbb{Z}$ denote the set of all the integers. For a real number $a$, denote by $N(a)$ the number of integers in the range $[1,a-1]$ not divisible by $q$, i.e., $N(a)=\left\lfloor a-1\right\rfloor-\left\lfloor(a-1)/q\right\rfloor$. Let $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/{3}\rfloor+1}]$ be integers. Let $k_{\delta}=\lfloor\log_{q}(\delta-1)\rfloor-h$. This implies that $q^{h+k_{\delta}}\leq\delta-1<q^{h+k_{\delta}+1}$. Let $\sum\limits_{\ell=0}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\delta-1$. If $k_{\delta}\geq m-2h$, let $s_{\delta}$ denote the smallest integer in $[m-2h-k_{\delta},k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$.
@@ IV The dimension of $\mathcal{C}_{(q,m,\delta)}$ for $2\leq\delta\leq q^{\lfloor({2m-1})/{3}\rfloor+1}$
$f(\delta)=\begin{cases}0,&\hbox{if }\delta\leq q^{h+1},\\ \left(\begin{aligned} &q^{2k_{\delta}-3}(k_{\delta}-1)(q-1)^{2}+N\left(\mu(\delta)+1\right)\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)\end{aligned}\right),&\hbox{if }\delta>q^{h+1},\end{cases}$ (27)
where $\mathcal{T}_{i}(\delta)=\left\{t\in\mathbb{Z}:q^{k_{\delta}-i}\leq t<\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h-i}\hbox{ and }q\nmid t\right\}$ and $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}.$
@@
$\widetilde{f}(\delta)=\begin{cases}0,&\hbox{if }\delta\leq q^{h},\\ \frac{1}{2}(\delta_{h}-1)\delta_{h}+N(\widetilde{\mu}(\delta)+1),&\hbox{if }q^{h}<\delta\leq q^{h+1},\\ \left(\begin{aligned} &(k_{\delta}-\frac{1}{2})q^{2k_{\delta}-2}(q-1)^{2}\\ &+\frac{1}{2}q^{k_{\delta}-1}(q-1)+N(\widetilde{\mu}(\delta)+1)\\ &+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)\end{aligned}\right),&\hbox{if }\delta>q^{h+1},\end{cases}$ (28)
@@
We also define the function $g(\delta)$ for integers $\delta\in[2,q^{\lfloor(2m-1)/{3}\rfloor+1}]$ as follow:
$g(\delta)=\begin{cases}0,&\hbox{if }\delta\leq q^{h},\\ \delta_{h}-1+\tau(\delta),&\hbox{if }q^{h}<\delta\leq q^{h+1},\\ N\left(\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\right)+\tau(\delta),&\hbox{if }\delta>q^{h+1}.\end{cases}$ (29)
@@
|$\left|\left\{(a_{m-1},\ldots,a_{0})\in Z_{q}^{m}:\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell}\leq t\hbox{ and }a_{0}>0\right\}\right|=N(t+1).$
@@
$\sum\limits_{t=q^{k}}^{q^{k+1}-1}N(t+1)=\begin{cases}\frac{1}{2}(q^{2}-q),&\hbox{if }k=0,\\ \frac{1}{2}q^{2k-1}(q-1)^{2}(q+1),&\hbox{if }k\geq 1.\\ \end{cases}$ (32)
@@
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i-1}+1)=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-3},&\hbox{if }i\in[-k+2,k-1],\\ \frac{1}{2}a(a-1)(q-1)q^{2k-2},&\hbox{if }i=k\hbox{ or }-k+1.\end{cases}$ (33)
@@
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i}+1)=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-2},&\hbox{if }i\in[-k+1,0)\cup(0,k-1],\\ \frac{1}{2}a(a-1)(q-1)q^{2k-1},&\hbox{if }i=k\hbox{ or }-k,\\ \frac{1}{2}(a-1)q^{k-1}(q-1)+\frac{1}{2}(a^{2}-1)q^{2k-2}(q-1)^{2},&\hbox{if }i=0.\end{cases}$ (34)
@@
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{S}\right|=\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1).$ (38)
@@
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|=\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)$ (39)
@@
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\right|=N(\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h})-N(q^{k_{\delta}}).$ (40)
@@
    $\left|\mathcal{A}_{k}(i)\right|=\begin{cases}\frac{1}{2}q^{2k-1}\left(q-1\right)^{2},&\hbox{if }i=k\hbox{ or }-k+1,\\ \frac{1}{2}q^{2k-3}(q-1)^{3}(q+1),&\hbox{if }i\in[-k+2,k-1].\end{cases}$ (41)
@@
    $\left|\mathcal{B}_{k}(i)\right|=\\ \begin{cases}\frac{1}{2}q^{2k}(q-1)^{2},&\hbox{if }i=k\hbox{ or }-k,\\ \frac{1}{2}q^{2k-2}(q-1)^{3}(q+1),&\hbox{if }i\in[-k+1,0)\cup(0,k-1].\end{cases}$ (42)
@@
$=$ $\sum\limits_{{k}=1}^{k_{\delta}-1}({k}-1)q^{2{k}-3}(q-1)^{3}(q+1)+\sum\limits_{{k}=1}^{k_{\delta}-1}q^{2{k}-1}(q-1)^{2}.$

### v3

@@ II Preliminaries
with $\mathcal{G}=\bigcup\limits_{a=b}^{b+\delta-2}C_{a}.$ It is clear that $\mathrm{deg}(g(x))=\left|\bigcup\limits_{a=b}^{b+\delta-2}C_{a}\right|$, where $\mathrm{deg}$ denotes the degree of a polynomial, and $|\cdot|$ denotes the size of a set. The dimension of $\mathcal{C}_{(q,m,\delta,b)}$ can be given by
@@
$\mathrm{dim}(\mathcal{C}_{(q,m,\delta)})=n-\sum\limits_{a\in\mathcal{L}(1,\delta)}\left|C_{a}\right|,$ (2)
where $\mathcal{L}(1,\delta)$ denotes the set of coset leaders in $1,\delta-1].$ Additionally, for $\delta^{{}^{\prime}}\geq\delta$, we can observe that
$\bigcup\limits_{a=1}^{\delta-1}C_{a}=\bigcup\limits_{a=1}^{\delta^{{}^{\prime}}-1}C_{a}$
if and only if any integer in $\delta,\delta^{{}^{\prime}}-1]$ is not a coset leader. Consequently, by recalling the definition of Bose distance $d_{B}$, we have
$d_{B}(\mathcal{C}_{(q,m,\delta)})=\delta^{{}^{\prime}},$ (3)
where $\delta^{{}^{\prime}}$ denotes the smallest coset leader in $\delta,n-1]$.
For simplicity, we use $a\mid b$ to denote that the integer $a$ divides the integer $b$, and $a\nmid b$ to denote that $a$ does not divide $b$. It is well known that a necessary condition for an integer $a$ being the coset leader of $C_{a}$ is $q\nmid a$. Therefore, we only need to focus on finding those integers that are neither divisible by $q$ nor a coset leader. We define $\mathcal{S}$ as the set of all such integers in [1,n1][1,n-1]. That is,
$\mathcal{S}$ $=$ $\{a\in[1,n-1]:q\nmid a\hbox{ and }a\hbox{ is not a coset leader}\}.$
@@
[47, Theorem 2.3] When $m$ is an odd integer, for any integer $\leq q^{h+1}, aa$ is the coset leader of $C_{a}$ if and only if $q\nmid a$. When $m$ is an even integer, for any integer $\leq 2q^{h}, aa$ is the coset leader of $C_{a}$ if and only if $q\nmid a$.
[…]
@@
When $m$ is an even integer, for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
@@
where $\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell}$ forms the $q$-adic expansion of the integer $a\in[0,n].$ Let $I_{m-1}$ be the identity matrix of order m1m-1, and let
@@
When $m$ is an even integer, we can use a similar argument as in the first paragraph to show that $l_{a}>\lfloor{m}/{3}\rfloor$ for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}).$ Note that $l_{a}$ divides $m$ and $m$ is even. It follows that $l_{a}={m}/{2}$ or $m$. Consequently, we have $l_{a}={m}/{2}=h$ if $aq^{h}\mathrm{\ mod\ }n=a$, and $l_{a}=m$ otherwise. Equivalently, the equality in (6) holds. This completes the proof. ∎
@@
Let $\mathcal{H}$ be the set of all integers $a\in[1,n-1]$ such that $a$ is a coset leader and $|C_{a}|={m}/{2}$. That is,
$\mathcal{H}=\left\{a\in[1,n-1]:a\hbox{ is a coset leader and }|C_{a}|=\frac{m}{2}\right\}.$
@@
$(\mathbf{0}_{h-k-1},a_{h+k}\ldots,a_{0})\\ =(a_{h-1},\ldots,a_{0},\mathbf{0}_{h-k-1},a_{h+k}\ldots,a_{h}).$ (10)
@@
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}\cdot q^{2i-1}$
if $i\in[1,k],$ and
$\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}\cdot q^{2i-1}$ $=$ $\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}$
$+\>\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}$
if $i\in[-k+1,0]$. Noticing that $0\leq\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}<1$ for $i\in[-k+1,0]$, and the two summations in (17) are both integers, we can conclude that (17) is equivalent to
@@
If $a\in\mathcal{A}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell},$ then $i$ is the smallest integer in [k+1,k][-k+1,k] such that $a_{h+i}>0$. Consequently, $\mathcal{A}_{k}(i)\cap\mathcal{A}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k+1,k]$.
Similarly, if $a\in\mathcal{B}_{k}(i)$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$, then $i$ is the smallest integer in [k,k][-k,k] such that $a_{h+i}>0$. As a result, $\mathcal{B}_{k}(i)\cap\mathcal{B}_{k}(j)=\varnothing$ for distinct integers $i,j\in[-k,k]$.
@@
$V(a)Q^{h-i+1}\\ =(\mathbf{0}_{h-k},a_{k+i-1},\ldots,a_{0},\mathbf{0}_{h-k},a_{h+k},\ldots,a_{h+i}).$ (23)
@@
With (13), it follows that
$(a_{k+i-1},\ldots,a_{2i-1})<(a_{h+k},\ldots,a_{h+i}).$
Consequently, we still have $V(a)Q^{h-i+1}<V(a)$, which implies $a\in\mathcal{S}$. By now, we have already demonstrated that $\bigcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i)\subseteq\mathcal{S}\cap[q^{h+k},q^{h+k+1}).$
Conversely, let us assume that $a\in\mathcal{S}\cap[q^{h+k},q^{h+k+1})$ is an integer with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$. We first have $a_{h+k}>0$. By the definition of $\mathcal{S}$ and Observation 3, we also have $a_{0}>0$ and $V(a)Q^{t}<V(a)$ for some integer $t\in[1,m-1]$. Noticing that $V(a)=(\mathbf{0}_{h-k},a_{h+k},\ldots,a_{0})$, it follows that the first hkh-k entries of $V(a)Q^{t}$ are all equal to zero. Considering $a_{h+k}>0$ and $a_{0}>0$, we have $h-k+1\leq t\leq h+k$. Consequently, we can conclude that
@@
Given that $a_{h+k}>0$, we can chose $i$ to be the smallest integer in [mht,k][m-h-t,k] such that $a_{h+i}>0$. Then $V(a)$ has the form specified in (14) with $(a_{k+i-1},\ldots,a_{0})=(\mathbf{0}_{h+t+i-m},a_{h+k-t},\ldots,a_{0})$. Furthermore, the inequality $h-k+1\leq t\leq h+k$ implies that $i\in[-k+1,k]$.
Additionally, if i>mhti>m-h-t, we can obtain (13) from the fact that $a_{h+k}>0$ and $a_{0}>0$. On the other hand, if i=mhti=m-h-t, then we can derive (13) from $V(a)Q^{t}<V(a)$ and $a_{0}>0$. Now we can conclude that $a\in\mathcal{A}_{k}(i)$ for some integer $i\in[-k+1,k]$. It follows that $\mathcal{S}\cap[q^{h+k},q^{h+k+1})\subseteq\bigcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i).$
We now can conclude from the above argument that $\bigcup\limits_{i=-k+1}^{k}\mathcal{A}_{k}(i)=\mathcal{S}\cap[q^{h+k},q^{h+k+1}).$ Finally, applying Remark 6 directly yields equation (21).
@@
[…]
Let $\mathbb{Z}$ denote the set of all the integers. For a real number $a$, denote by $N(a)$ the number of integers in the range [1,a1][1,a-1] not divisible by $q$, i.e., $N(a)=\left\lfloor a-1\right\rfloor-\left\lfloor(a-1)/q\right\rfloor$. Let $m\geq 4$ and $\delta\in[2,q^{\lfloor(2m-1)/{3}\rfloor+1}]$ be integers. Let $k_{\delta}=\lfloor\log_{q}(\delta-1)\rfloor-h$. This implies that $q^{h+k_{\delta}}\leq\delta-1<q^{h+k_{\delta}+1}$. Let $\sum\limits_{\ell=0}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\delta-1$. If $k_{\delta}\geq m-2h$, let $s_{\delta}$ denote the smallest integer in $m-2h-k_{\delta},k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$.
@@
$f(\delta)=\begin{cases}0,&\delta\leq q^{h+1},\\ \left(\begin{aligned} &q^{2k_{\delta}-3}(k_{\delta}-1)(q-1)^{2}\\ &+N\left(\mu(\delta)+1\right)\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1)\end{aligned}\right),&\delta>q^{h+1},\end{cases}$ (25)
where $\mathcal{T}_{i}(\delta)=\left\{t\in\mathbb{Z}:q^{k_{\delta}-i}\leq t<\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h-i},q\nmid t\right\}$ and $\mu(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta}+1)}\right\}.$
@@
$\widetilde{f}(\delta)=\begin{cases}0,\hskip-10.0pt&\hskip-6.0pt\delta\leq q^{h},\\ \frac{1}{2}(\delta_{h}-1)\delta_{h}+N(\widetilde{\mu}(\delta)+1),\hskip-10.0pt&\hskip-6.0ptq^{h}<\delta\leq q^{h+1},\\ \left(\begin{aligned} &(k_{\delta}-\textstyle\frac{1}{2})q^{2k_{\delta}-2}(q-1)^{2}\\ &+\textstyle\frac{1}{2}q^{k_{\delta}-1}(q-1)\\ &+N(\widetilde{\mu}(\delta)+1)\\ &+\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)\end{aligned}\right),\hskip-10.0pt&\hskip-6.0pt\delta>q^{h+1},\end{cases}$ (26)
@@
In addition, we define the function $g(\delta)$ for integers $\delta\in[2,q^{\lfloor(2m-1)/{3}\rfloor+1}]$ as follow:
$g(\delta)=\begin{cases}0,&\delta\leq q^{h},\\ \delta_{h}-1+\tau(\delta),&q^{h}<\delta\leq q^{h+1},\\ N\left(\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\right)+\tau(\delta),&\delta>q^{h+1}.\end{cases}$ (27)
@@
$\left|\left\{(a_{m-1},\ldots,a_{0})\in Z_{q}^{m}:\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell}\leq t,a_{0}>0\right\}\right|\\ =N(t+1).$ (30)
@@
[…]
@@
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i-1}+1)$ $=$ $\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-3},\\ \quad\hbox{if }i\in[-k+2,k-1],\\ \frac{1}{2}a(a-1)(q-1)q^{2k-2},\\ \quad\hbox{if }i=k\hbox{ or }-k+1,\end{cases}$ (31)
@@
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}}N(tq^{2i}+1)$
$\quad=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-2},\\ \quad\hbox{if }i\in[-k+1,k-1]\setminus\{0\},\\ \frac{1}{2}a(a-1)(q-1)q^{2k-1},\\ \quad\hbox{if }i=k\hbox{ or }-k,\\ \frac{1}{2}(a^{2}-1)(q-1)^{2}q^{2k-2}+\frac{1}{2}(a-1)(q-1)q^{k-1},\\ \quad\hbox{if }i=0.\end{cases}$ (32)
@@
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{S}\right|\\ =\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}N(tq^{2i-1}+1).$ (36)
@@
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h-k_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap(\mathcal{S}\cup\mathcal{H})\right|\\ =\sum\limits_{i=-k_{\delta}}^{k_{\delta}}\sum\limits_{t\in{\mathcal{T}}_{i}(\delta)}N(tq^{2i}+1)$ (37)
@@
    $\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\right|\\ =N\left(\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\right)-N\left(q^{k_{\delta}}\right).$ (38)
@@
    $\left|\mathcal{A}_{k}(i)\right|=\begin{cases}\frac{1}{2}q^{2k-1}\left(q-1\right)^{2},\\ \quad\hbox{if }i=k\hbox{ or }-k+1,\\ \frac{1}{2}q^{2k-3}(q-1)^{3}(q+1),\\ \quad\hbox{if }i\in[-k+2,k-1].\end{cases}$ (39)
@@
    $\left|\mathcal{B}_{k}(i)\right|=\\ \begin{cases}\frac{1}{2}q^{2k}(q-1)^{2},\\ \quad\hbox{if }i=k\hbox{ or }-k,\\ \frac{1}{2}q^{2k-2}(q-1)^{3}(q+1),\\ \quad\hbox{if }i\in[-k+1,k-1]\setminus\{0\}.\end{cases}$ (40)
@@
$=$ $\sum\limits_{{k}=1}^{k_{\delta}-1}({k}-1)q^{2{k}-3}(q-1)^{3}(q+1)$
$+\>\sum\limits_{{k}=1}^{k_{\delta}-1}q^{2{k}-1}(q-1)^{2}.$